OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3? (0, 3) (1, -1) (-1, 4)

OpenStudy (anonymous):

(0, 3) (1, -1) (-1, 4) Plug in the x and y values and simplify to see if the statement is true (3)=2(0) -3 3 = 0 - 3 3= -3 no -1 = 2(1) - 3 -1 = 2 - 3 -1 = - 1 yes 4 = 2 (-1) - 3 4 = -2 -3 4 = -5 no

OpenStudy (anonymous):

what about this one? Which of the following points does not lie on the line whose equation is x+ y = 7? (-10, 17) (13/3, 8/3) (-1, -6)

OpenStudy (anonymous):

Do the same thing, plug in the x and y values and simplify to see it the statement is true -10+17 = 7 7 = 7 yes 13/3 + 8/3 = 7 7 = 7 yes -1 -6 = 7 -7 = 7 no

OpenStudy (anonymous):

okay thanks alot